# Search for spatial coincidence between IceCube neutrinos and gamma-ray bright red dwarfs

### v1

@@ I Introduction
In this work, we look for high energy neutrinos from red dwarfs. Red dwarfs (sometimes also referred to as M-dwarfs) are one of the smallest stars on the main sequence with masses between 0.075-0.5 $M_{\odot}$, having surface temperatures between 2,500-5,000 K Edgeworth (1946), and frequently emit flares with energies between $10^{32}$ and $10^{35}$ ergs with a power-law index between 1.7-2.4 Yang et al. (2017). Energetic outbursts have been detected from red dwarfs at high energies from hard X-rays to TeV gamma-rays Osten et al. (2016); Song and Paglione (2020); Sinitsyna et al. (2019). Most recently, the SHALON TeV gamma-ray telescope detected very high energy gamma-ray emission between 800 GeV and 20 TeV from eight red dwarfs, namely, V388 Cas, V547 Cas, V780 Tau, V962 Tau, V1589 Cyg, GJ 1078, GJ 3684 and GL 851.1 Sinitsyna et al. (2019). Subsequently, these red dwarfs have been proposed as sources of ultra high energy cosmic rays Sinitsyna et al. (2022). However, no corresponding gamma-ray emission was seen at GeV Huang et al. (2024) or MeV energies Shrivastava et al. (2024)Song and Paglione (2020) reported a $5\sigma$ detection of pulsed GeV $\gamma$-ray emission from another M-dwarf TVLM 513-46546 with Fermi-LAT. A very hard X-ray outburst was detected by the SWIFT-BAT telescope from another M-dwarf, viz DG CVn from 0.3-100 keV in April 2014 Osten et al. (2016), although no coincident gamma-ray emission was seen from Fermi-LAT Loh et al. (2017). Although there are no estimates for high energy neutrino flux emission from red dwarfs, hadronic models for gamma-ray emission have been proposed Shrivastava et al. (2024). One would also then expect associated neutrino emission along with the high energy gamma-rays.
@@ II Dataset
The neutrinos used in this work are taken from the IceCube 10 year muon track data Abbasi et al. (2021). IceCube is a detector located at the South Pole which detects Cherenkov photons emitted by secondary charged particles produced in neutral current (NC) or charged current (CC) interactions, when neutrinos interact with 2800-thick glacial ice. This dataset consists of 1,134,431 neutrinos, which were collected between April 2008 (IC-40) and July 2018 (IC 86-V11) from four different phases of the experiment, each having different livetime. Note that for our analysis we have used the augmented dataset analyzed in  Zhou and Beacom (2022) [^1] This dataset is available at https://github.com/beizhouphys/IceCube_data_2008--2018_double_counting_corrected, which corrects for 38 duplicate events found in the IceCube public muon track data due to angular reconstruction error Zhou and Beacom (2022). For each neutrino, the dataset consists of right ascension (RA), declination ($\delta$), error in position and reconstructed muon energy.
@@ III Analysis and Results
where $|\theta_{i}-\theta_{s}|$ is the angular distance between the neutrino and the red dwarf, $\sigma_{i}$ is the angular uncertainty in the neutrino position, in radians.The background PDF can be obtained from the solid angle within the declination ($\delta$) of ±$\pm 5^{\circ}$ around each red dwarf and is given as follows:
@@
Now to ascertain the significance of the signal, we define a detection statistic ($TS$) as follows:
$TS(n_{s})=2\log\frac{\mathcal{L}(n_{s})}{\mathcal{L}(0)}$ (4)
For the null hypothesis, $TS(n_{s})$ behaves like a $\chi^{2}$ distribution for one degree of freedom Wilks (1938). The detection significance can be obtained from $\sqrt{TS}$. This statistics is also used in gamma-ray astronomy (eg. Manna and Desai (2024)). For a statistically significant detection of $>5\sigma$, TS $>25$. For each of our red dwarf, we calculate the best-fit $n_{s}$ which maximizes $TS$ according to Eq. 4. These values of TS$(n_{s})$, along with the best-fit values of $n_{s}$ can be found in Table 1. As we can see, none of the red dwarfs show a TS value $>25$. The largest TS value is seen for GJ 1078 with TS of 2.38, which is less than $2\sigma$. Therefore, the observed signal events are consistent with background and there is no evidence for any spatial association between red dwarfs and IceCube neutrinos, which are observed as muon tracks.
@@ IV Conclusions
In this work, we searched for a spatial coincidence between IceCube neutrinos and 10 gamma-ray bright red dwarfs. The neutrinos used in this work are taken from the publicly released 10 year muon track data observed between 2008-2018, and consisting of 113443 neutrinos. Eight of the 10 red dwarfs were detected at TeV gamma rays by the SHALON telescope. Two others were detected in gamma-rays and X-rays from Fermi-LAT and SWIFT. We analyzed these data using the unbinned maximum likelihood method. Our results for the detection significance ($TS(n_{s})$) for all the red dwarfs can be found in Table 1. All the red dwarfs show a TS value less than two. Therefore, we conclude that none of the above red dwarfs contribute to the Ice Cube diffuse neutrino flux.
Red dwarf RA () $\delta(^{\circ}$) $n_{s}$ $TS_{max}$
V388 Cas 15.8326 +62.365 0.0 0.0
V547 Cas 8.1229 +67.235 0.0 0.0
V780 Tau 85.107 +24.802 6.00 0.35
V962 Tau 86.466 +22.879 0.0 0.0
V1589 Cyg 310.705 +41.38 0.0 0.0
GJ 1078 80.95 +22.54 20.39 2.38
GJ 3684 176.77 +70.03 0.0 0.0
GL 851.1 333.0267 +31.56 0.0 0.0
TVLM 513-46546 225.284 22.834 0.0 0.0
DG CVn 202.944 29.28 2.42 0.05
Table 1: Results from spatial coincidence analysis between IceCube neutrinos and red dwarfs. None of the red dwarfs show a detection significance $>5\sigma$, corresponding to TS$>25$. Therefore, we do not find any spatial association between the red dwarfs and IceCube neutrinos.
@@ Acknowledgments
We are grateful to Vibhavasu Pasumarti for making the analysis codes used in  Pasumarti and Desai (2022) publicly available.

### v2

@@ I Introduction
In this work, we look for high energy neutrinos from red dwarfs. Red dwarfs (sometimes also referred to as M-dwarfs) are one of the smallest stars on the main sequence with masses between 0.075-0.5 $M_{\odot}$, having surface temperatures between 2,500-5,000 K Edgeworth (1946), and frequently emit flares with energies between $10^{32}$ and $10^{35}$ ergs with a power-law index between 1.7-2.4 Yang et al. (2017). Energetic outbursts have been detected from red dwarfs at high energies from hard X-rays to TeV gamma-rays Osten et al. (2016); Song and Paglione (2020); Sinitsyna et al. (2019). Most recently, the SHALON TeV gamma-ray telescope detected very high energy gamma-ray emission between 800 GeV and 20 TeV from eight red dwarfs, namely, V388 Cas, V547 Cas, V780 Tau, V962 Tau, V1589 Cyg, GJ 1078, GJ 3684 and GL 851.1 Sinitsyna et al. (2019). Subsequently, these red dwarfs have been proposed as sources of ultra high energy cosmic rays Sinitsyna et al. (2022, 2021). However, no corresponding gamma-ray emission was seen at GeV Huang et al. (2024) or MeV energies Shrivastava et al. (2024)Song and Paglione (2020) reported a $5\sigma$ detection of pulsed GeV $\gamma$-ray emission from another M-dwarf TVLM 513-46546 with Fermi-LAT. A very hard X-ray outburst was detected by the SWIFT-BAT telescope from another M-dwarf, viz DG CVn from 0.3-100 keV in April 2014 Osten et al. (2016), although no coincident gamma-ray emission was seen from Fermi-LAT Loh et al. (2017). Although there are no estimates for high energy neutrino flux emission from red dwarfs, these red dwarfs could generate cosmic rays with energies up to $10^{14}$ eV in their flaring states Sinitsyna et al. (2021). TeV gamma-rays detected from these red dwarfs indicate that these red dwarfs produce flares with total energies of $10^{32}-10^{35}$ ergs Sinitsyna et al. (2022, 2021). If one considers the statistics of M-dwarfs in our galaxy and assuming the frequency of flares to be around 36/year for about 10% of the observed red dwarfs, the total energy released from these flares matches the total energy of cosmic rays in the galactic disk of $\sim 10^{54}$ ergs Sinitysna et al. (2019). During these flares the accelerated protons with energies up to $10^{14}$ eV undergo spallation with the atmosphere outside the star Sinitsyna et al. (2021). Note that similar mechanisms for production of TeV gammma-rays have also been proposed for the Sun Zhou et al. (2017). As argued in  Sinitsyna et al. (2021), these interactions produce pions. Consequently, one would also expect neutrino emission from the aforementioned hadronic interactions along with gamma-rays produced from pion decay. Therefore, these considerations provide enough motivation to search for neutrino emissions to complement recent searches for MeV Shrivastava et al. (2024) and GeV Huang et al. (2024) gamma-rays from these red dwarfs.
@@ II Dataset
The neutrinos used in this work are taken from the IceCube 10 year muon track data Abbasi et al. (2021). IceCube is a detector located at the South Pole, which detects neutrinos through the Cherenkov radiation emitted by the charged leptons traveling through the ice.
This dataset consists of 1,134,431 track-like neutrino candidates, which were collected between April 2008 (IC-40) and July 2018 (IC 86-V11) from four different phases of the experiment, each having different livetime. Note that for our analysis we have used the augmented dataset analyzed in  Zhou and Beacom (2022) [^1] This dataset is available at https://github.com/beizhouphys/IceCube_data_2008--2018_double_counting_corrected, which corrects for 38 duplicate events found in the IceCube public muon track data due to angular reconstruction error Zhou and Beacom (2022). For each neutrino, the dataset consists of right ascension (RA), declination ($\delta$), error in the position, and reconstructed muon energy.
@@ III Analysis and Results
where $|\theta_{i}-\theta_{s}|$ is the angular distance between the neutrino and the red dwarf, $\sigma_{i}$ is the angular uncertainty in the neutrino position, in radians.The background PDF can be obtained from the solid angle within the declination ($\delta$) of ±$\pm 5^{\circ}$ around each red dwarf and is given as follows Braun et al. (2008):
@@
We then evaluate $\mathcal{L}(n_{s})$ for different values of $n_{s}$, and designate the value of $n_{s}$, for which this likelihood is maximized as $\hat{n}_{s}$. Now to ascertain the significance of the signal, we define the test statistics ($TS$) based on the likelihood ratio, with respect to the null hypothesis in the absence of a signal, ($\mathcal{L}(0)$) as follows:
$TS=2\log\frac{\mathcal{L}(\hat{n}_{s})}{\mathcal{L}(0)}.$ (4)
For the null hypothesis, $TS$ behaves like a $\chi^{2}$ distribution for one degree of freedom Wilks (1938). The detection significance can be obtained from $\sqrt{TS}$. This statistics is also used in gamma-ray astronomy (eg. Manna and Desai (2024)). For a statistically significant detection of $>5\sigma$, TS $>25$. For each of our red dwarfs, we determine $\hat{n}_{s}$ and evaluate $TS$ according to Eq 4. These values of TS, along with the values of $\hat{n}_{s}$ can be found in Table 1. As we can see, none of the red dwarfs show a TS value $>25$. The largest TS value is seen for GJ 1078 with TS of 2.38, whose significance is less than $2\sigma$. Therefore, the observed signal events are consistent with background, and there is no evidence for any spatial association between red dwarfs and IceCube neutrinos, which are observed as muon tracks. Consequently, we calculate the 95% c.l. upper flux limits by calculating the value of $n_{s}$ for which $\Delta(TS)=-3.84$ Zyla et al. (2020), where $\Delta TS=TS-TS_{max}$.
@@ IV Conclusions
In this work, we searched for a spatial coincidence between IceCube neutrinos and 10 gamma-ray bright red dwarfs. The neutrinos used in this work are taken from the publicly released 10 year muon track data observed between 2008-2018, and consisting of 113443 neutrinos. Eight of the 10 red dwarfs were detected at TeV gamma rays by the SHALON telescope. Two others were detected in gamma-rays and X-rays from Fermi-LAT and SWIFT. We analyzed these data using the unbinned maximum likelihood method. A tabular summary of our results for all the red dwarfs can be found in Table 1. All the red dwarfs show a test statistic value of less than two, with significance less than $2\sigma$. Therefore, we conclude that none of the above red dwarfs contribute to the Ice Cube diffuse neutrino flux.
Red dwarf RA () $\delta(^{\circ}$) $\hat{n}_{s}$ $TS$ Upper limit
V388 Cas 15.8326 +62.365 0.0 0.0 16.2
V547 Cas 8.1229 +67.235 0.0 0.0 16.7
V780 Tau 85.107 +24.802 6.00 0.35 29.8
V962 Tau 86.466 +22.879 0.0 0.0 13.8
V1589 Cyg 310.705 +41.38 0.0 0.0 12.7
GJ 1078 80.95 +22.54 20.39 2.38 51.2
GJ 3684 176.77 +70.03 0.0 0.0 14.9
GL 851.1 333.0267 +31.56 0.0 0.0 11.1
TVLM 513-46546 225.284 22.834 0.0 0.0 20.8
DG CVn 202.944 29.28 2.42 0.05 27.2
Table 1: Results from spatial coincidence analysis between IceCube neutrinos and red dwarfs. None of the red dwarfs show a detection significance $>5\sigma$, corresponding to TS$>25$. Therefore, we do not find any spatial association between the red dwarfs and IceCube neutrinos. The last column shows the 95% c.l. upper limit on the number of signal neutrino events.
@@ Acknowledgments
We are grateful to Vibhavasu Pasumarti for making the analysis codes used in  Pasumarti and Desai (2022) publicly available and the anonymous referee for constructive feedback on the manuscript.